An acid spill occurred at a metal processing plant in which a vat of hydrochloric acid ruptured spilling 30 000 L into the surrounding soils. The acid from a soil sample was extracted and analyzed by titration. A 20.0 mL sample required 16.0 mol/L 2.00 mol.L calcium hydroxide solution to undergo neutralization.
1
Expert's answer
2019-08-12T07:58:13-0400

n(Ca(OH)2) = C(Ca(OH)2) × V(Ca(OH)2) = 2.00 mol/L × 0.016 L = 0.032 mol.


0.032 mol   X mol

 Ca(OH)2 + 2HCl = CaCl2 + 2H2O

   1 mol      2 mol


n(HCl) = X = 0.064 mol.

C(HCl) = n(HCl) / V(HCl) = 0.064 mol / 0.020 L = 3.2 mol/L.